# Fourier123: One Image to High-Quality 3D Object Generation with Hybrid Fourier Score Distillation

### v1

@@ Abstract
Single image-to-3D generation is pivotal for crafting controllable 3D assets. Given its underconstrained nature, we leverage geometric priors from a 3D novel view generation diffusion model and appearance priors from a 2D image generation method to guide the optimization process. We note that a disparity exists between the training datasets of 2D and 3D diffusion models, leading to their outputs showing marked differences in appearance. Specifically, 2D models tend to deliver more detailed visuals, whereas 3D models produce consistent yet over-smooth results across different views. Hence, we optimize a set of 3D Gaussians using 3D priors in spatial domain to ensure geometric consistency, while exploiting 2D priors in the frequency domain through Fourier transform for higher visual quality. This 2D-3D hybrid Fourier Score Distillation objective function (dubbed hy-FSD), can be integrated into existing 3D generation methods, yielding significant performance improvements. With this technique, we further develop an image-to-3D generation pipeline to create high-quality 3D objects within one minute, named Fourier123. Extensive experiments demonstrate that Fourier123 excels in efficient generation with rapid convergence speed and visual-friendly generation results.
@@ 1 Introduction
One image-to-3D generation is the process of producing exquisite high-fidelity 3D assets from a given image, which offers substantial advantages for empowering non-professional users to engage in 3D asset creation. However, due to the lack of constraints, this task remains challenging despite decades of development [1; 2; 3]. Recent advances in deep learning-based generative models [4; 5] have inspired an increasing number of 3D generation methods. Among these, three strategies are widely studied and have achieved State-Of-The-Art (SOTA) results: 1) optimization-based 2D lifting methods, 2) novel view generation diffusion models, and 3) 3D native methods.
Considering the impressive advancements in image generation [5; 6], Dreamfusion [7] proposed Score Distillation Sampling (SDS) to distill 3D geometry and appearance from powerful 2D diffusion models [8], inspiring the development of 2D lifting methods. Early work [9; 10] tended to use Neural Radiance Fields (NeRF) to model 3D, which required hours of training. Although recent studies [11; 12] adopt 3D Gaussian Splatting (3DGS) to accelerate generation, they still require a few minutes for optimization. On the other hand, some researchers fine-tuned existing 2D image diffusion models on 3D datasets, enabling these models to produce novel views around an object [13; 14; 15]. Whilst 3D native methods directly generate 3D-consistent assets within seconds [16; 17], e.g., mesh, point clouds, or NeRF, etc., they still suffer from limited visual quality and realism. To this end, we propose Fourier123, an efficient image-to-3D generation pipeline that generates visually appealing 3D assets within one minute. As shown in Fig. 1, our method predicts lifelike content for occluded regions while maintaining consistency with the input image.
@@
Figure 2: Frequency analysis of Stable Diffusion (SD) and Zero-1-to-3 (Zero123). Discrete Fourier Transform (DFT) converts the results of SD and Zero123 to frequency domain. In the upper row, image produced by SD exhibits high visual quality but distorts identity. In the lower row, Zero123 generates an image that matches the input but is over-smooth. Their frequency results, i.e., 3DF of Zero123 and 2DF of SD, are also different. Frequency difference showcases the absolute value of their amplitude subtraction. Here we omit conditions of diffusion models for concise presentation, and all details can be found in Fig. 3.
Motivation.
Recent optimization-based 3D generation studies [18; 19] have combined 2D and 3D diffusion models to generate 3D assets. These methods calculate pixel-level losses with 2D and 3D models respectively, and roughly combine them to train a 3D representation, but ignore the differences in generated images caused by their distinct training data domains. As shown in Fig. 2, Zero-1-to-3 (Zero123) and Stable Diffusion (SD) tend to produce different reference images, directly combining them in the spatial domain may confuse training. Intuitively, SD produces a higher quality image but distorts identity, while Zero123 faithfully generates a view but with blurred appearance. By observing results of SD (2DF) and Zero123 (3DF) in the frequency domain through Fourier transform, we found their amplitude components are different, especially at the intermediate frequencies. We believe the frequency results of SD are more consistent with the frequency distribution of high-quality images, whereas the spatial domain results of Zero123 are aligned with the identity of the input image.
Thus, different from previous methods that merely combine the supervision of 2D and 3D diffusion models in the spatial domain, we, for the first time, supervise the distillation process using a 3D diffusion model in the spatial domain and a 2D diffusion model in the frequency domain, fully utilizing the geometric priors of the former and the high-quality generation capabilities of the latter. This 2D-3D hybrid Fourier Score Distillation is called hy-FSD, which brings performance gains to existing optimization-based 3D generation methods. Building on hy-FSD, we further develop Fourier123, an efficient image-to-3D pipeline that generates a 3D object within one minute. To accelerate training, we adopt 3DGS for its fast convergence properties. Meanwhile, we employ the recently proposed Large Gaussian Model (LGM) [20] to initialize Gaussians, similar to the initialization in previous optimization-based methods [12; 18].
@@
    We pioneer the exploration of an optimization-based 3D generation that utilizes both spatial and frequency domains. We propose the 2D-3D hybrid Fourier Score Distillation (hy-FSD), significantly enhancing the performance of existing 3D generation methods.
@@
    We develop an efficient image-to-3D generation pipeline, Fourier123, utilizing hy-FSD. It enables high-quality 3D generation within one minute on a single NVIDIA 4090 GPU.
@@
    Extensive experiments confirm that Fourier123 effectively produces 3D assets with reliable structure and elegant appearance, fully integrating the generation capabilities of the image diffusion model and the 3D priors of the novel view generation model.
@@ 2 Related Work
Recently, various 3D representation techniques have been proposed for a range of 3D tasks. Wang et al. [21] employed volumetric rendering and reconstructed object surfaces by training an implicit network. Mildenhall et al. [22] further proposed NeRF, an end-to-end model popular for enabling 3D optimization with only 2D supervision. NeRF has inspired numerous subsequent studies, including 3D reconstruction [23; 24; 25; 26; 27] and generation [7; 10; 28; 29], but it consumes excessive time for optimization due to its computationally expensive forward and backward passes. Although some methods [30; 31; 32] attempted to accelerate training, the recent developed 3D Gaussian Splatting (3DGS) [33] achieved real-time rendering with faster training speed, and is considered a viable alternative 3D representation to NeRF. Its efficient differentiable splatting mechanism and representation design enable fast convergence and faithful reconstruction [34; 35; 36]. Recent studies on 3D generation [11; 12] have adopted 3DGS to achieve faster and higher quality generation. In this work, we also employ 3DGS as the representation technique and make the first attempt to realize the optimization-based methods in both spatial and frequency domains, improving generation quality.
@@ 2.2 Image-to-3D Generation
Image-to-3D generation aims to create 3D assets from a single reference image. This problem is also known as single-view 3D reconstruction, but such reconstruction settings [37; 38; 39] are limited by uncertainty modeling and often produce blurry results. Recently, diffusion models [5; 40] have achieved notable success in image generation, including text-to-image (T2I) [8; 41] and novel view synthesis [13; 15; 17; 42; 43]. Several methods have attempted to extend 2D image models for 3D generation [7; 10; 29; 44; 45], but they suffer from long optimization times as they require frequent generation of 2D images to train 3D representations. To address this, some studies explicitly injected camera parameters into 2D diffusion models for zero-shot novel view synthesis [13; 42]. Other methods have tried to build end-to-end large reconstruction models to generate 3D assets with a single forward process [16; 20; 46; 47]. However, the quality of the 3D assets generated by these two methods remains rough. Qian et al. [19] found that the T2I model [6] has impressive 2D generation capabilities, while Zero-1-to-3 [13] tends to generate reliable 3D structures. They used both 2D and 3D priors to generate 3D. In this paper, we combine 2D and 3D priors from the spatial and frequency domains, respectively, fully utilizing their respective advantages and enhancing 3D generation quality.
@@ 3 Preliminary
We use 3DGS [33] as our 3D representation technique. 3DGS uses anisotropic Gaussians to represent scenes, defined by a center position $\mathbf{\mu}\in\mathbb{R}^{3}$ and a covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$, which is decomposed into a scaling factor $\mathbf{s}\in\mathbb{R}^{3}$ and a rotation factor $\mathbf{r}\in\mathbb{R}^{4}$. Additionally, the color of each 3D Gaussian is defined by spherical harmonic (SH) coefficients $\mathbf{h}\in\mathbb{R}^{3\times(k+1)^{2}}$ for order $k$, along with an opacity value $\sigma\in\mathbb{R}$. The 3D Gaussian can be queried as:
@@
where $\mathbf{x}$ represents the position of the query point. To compute the color of each pixel, it uses a typical neural point-based rendering [48; 49]. Let $\mathbf{C}\in\mathbb{R}^{H\times W\times 3}$ represent the color of rendered image, where $H$ and $W$ are the height and width respectively. The rendering process is outlined as:
@@
where $\alpha_{t}\in(0,1)$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The denoiser $\bm{\epsilon}_{\theta}$ is trained by minimizing the noise reconstruction loss conditioned on $\mathbf{y}$ from pre-trained language models like CLIP [50]:
@@
where $\mathcal{C}(\mathbf{R},\mathbf{T})$ is the embedding of the input view and relative camera extrinsics. Once the model is trained, the user can input the image $\mathbf{x}$ and the external parameters $\mathbf{R},\mathbf{T}$ of the camera to obtain the target image under the corresponding pose.
@@
Score Distillation Sampling (SDS) [7; 29; 44] utilizes pre-trained text-to-image diffusion models for text-conditioned 3D generation. Specifically, SDS optimizes the parameters $\bm{\phi}$ of a differentiable 3D representation, such as a neural radiance field or 3DGS, by leveraging the gradient of the loss $\mathcal{L}_{\textrm{SDS}}$ with respect to $\bm{\phi}$:
@@
where $\mathbf{x}=g(\bm{\phi},c)$ represents an image rendered from the 3D representation $\bm{\phi}$ by the renderer $g$ under a specific camera pose $c$. The weighting function $w(t)$ depends on the timestep $t$, and the noise $\bm{\epsilon}$ is added to $\mathbf{z}=\mathcal{E}({\mathbf{x}})$ following Eq. 3 at timestep $t$. Its key insight is to enforce the rendered image of the learnable 3D representation to adhere to the distribution of the pre-trained diffusion model. However, relying on the 2D prior alone does not adequately capture prolific 3D geometry. Recent work, Magic123 [19], combines 3D priors of a novel view diffusion model to further improve generation quality. But it still remains two key drawbacks. First, this NeRF-based method requires numerous time for textual inversion and optimization. Second, it ignores the appearance conflict between the results of SD and Zero123 (as shown in Fig. 2), resulting in suboptimal results. In contrast, we use 3DGS to acclerate training, and convert the results of SD to the frequency domain for indirect supervision of appearance, without the requirement of textual inversion. The results of Zero123 are also combined in the spatial domain to optimize object structure.
@@ 4 Proposed Fourier123
We present Fourier123, an efficient image-to-3D generation method. A novel score distillation sampling function is developed by supervising in both spatial and frequency domains, termed hy-FSD (hybrid Fourier Score Distillation). This enables the produced 3D assets to benefit from the high-quality generation capability of a text-to-image (T2I) model (i.e., 2D model) and the geometric prior of a novel-view generation diffusion model (i.e., 3D model) simultaneously. Experiments validate that hy-FSD can be applied to existing 3D generation baselines to improve their performance. With hy-FSD, Fourier123 generates more reliable geometric structures and high-quality appearances.
@@ 4.1 Hybrid Fourier Score Distillation
The Discrete Fourier Transform (DFT) has been extensively employed in analyzing the frequency components of images. In the case of multichannel color images, the DFT $\mathcal{F}$ is computed and applied independently to each channel. For simplicity, the notation pertaining to channels is omitted in the presented formulas. For an image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, $\mathcal{F}$ converts it to the frequency domain as the complex component $\mathbf{X}$, expressed as:
$\mathcal{F}(\mathbf{x})_{u,v}=\mathbf{X}_{u,v}=\frac{1}{\sqrt{HW}}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\mathbf{x}_{h,w}e^{-j2\pi(u\frac{h}{H}+v\frac{w}{W})}.$ (7)
This function can be efficiently implemented with the FFT algorithm described in [51]. Note that its amplitude component $\mathcal{A}(\mathbf{x})_{u,v}$ can be formulated as:

### v2

# Hybrid Fourier Score Distillation for Efficient One Image to 3D Object Generation

@@ Abstract
Single image-to-3D generation is pivotal for crafting controllable 3D assets. Given its under-constrained nature, we attempt to leverage 3D geometric priors from a novel view diffusion model and 2D appearance priors from an image generation model to guide the optimization process. We note that there is a disparity between the generation priors of these two diffusion models, leading to their different appearance outputs. Specifically, image generation models tend to deliver more detailed visuals, whereas novel view models produce consistent yet over-smooth results across different views. Directly combining them leads to suboptimal effects due to their appearance conflicts. Hence, we propose a 2D-3D hybrid Fourier Score Distillation objective function, hy-FSD. It optimizes 3D Gaussians using 3D priors in spatial domain to ensure geometric consistency, while exploiting 2D priors in the frequency domain through Fourier transform for better visual quality. hy-FSD can be integrated into existing 3D generation methods and produce significant performance gains. With this technique, we further develop an image-to-3D generation pipeline to create high-quality 3D objects within one minute, named Fourier123. Extensive experiments demonstrate that Fourier123 excels in efficient generation with rapid convergence speed and visually-friendly generation results.
@@ 1 Introduction
One image-to-3D generation is the process of producing exquisite high-fidelity 3D assets from a given image, which offers substantial advantages for empowering nonprofessional users to engage in 3D asset creation. However, due to the lack of constraints, this task remains challenging despite decades of development [1; 2; 3]. Recent advances in deep learning-based generative models [4; 5] have inspired an increasing number of 3D generation methods and achieved State-Of-The-Art (SOTA) effects [6; 7], which are mainly divided into three strategies: 1) optimization-based 2D lifting methods, 2) novel view generation diffusion models, and 3) 3D native methods.
As existing image generation methods [5; 8] have been able to produce exquisite images, optimization-based methods attempt to using powerful 2D image generation models to achieve 3D generation [9; 7]. However, on the one hand, these optimization methods require tedious time for training. On the other hand, existing optimization strategy still cannot fully ustilizing the generation ability of 2D diffusion models in 3D generation, and remain some inherent problems such as Janus problem, leading to limited quality 3D generation. More recently, novel view diffusion models [10; 11; 12] and 3D native models [13; 6] that directly generate 3D objects are proposed to generate multi-view images or 3D assets within seconds. However, they are mainly trained or fine-tuned on 3D object datasets, which makes their generation ability relatively weak compared with image generation models, restricting the quality of 3D generation. To this end, we expect to develop an efficient score distillation function so that 3D generation can fully benefit from the leading generation capabilities of image generation methods for better effects.
@@
Figure 2: Frequency analysis of Stable Diffusion (SD) and Zero-1-to-3 (Zero123). Discrete Fourier Transform (DFT) converts the results of SD ($S_{2D}$) and Zero123 ($S_{3D}$) to frequency domain and here we visualize their amplitude components. In the upper row, $S_{2D}$ exhibits high visual quality but distorts content structure. In the lower row, $S_{3D}$ matches the input but is over-smooth. Their frequency amplitudes, $F_{3D}$ and $F_{2D}$, are also different. We train with $S_{3D}$ for its fidelity, and $F_{2D}$ for finer details. More details can be found in Fig. 3.
Motivation. (1) 2D SDS-based 3D generation. With the powerful generation capability of 2D Stable Diffusion (SD) [8], DreamFusion [14] utilizes SD to generate pseudo-GT views for training NeRF to achieve text-to-3D. However, SD is a 2D generative model that struggles to ensure good multi-view consistency, so a high Classifier-Free Guidance (CFG) value [15] is necessary but it decreases generation quality. Additionally, SD tends to generate forward-facing images [10], which is known as the Janus problem. These limitations restrict the ability to use SD for 3D generation directly. (2) 3D SDS-based 3D generation. Zero123 [10] fine-tunes the pre-trained 2D SD on 3D object-level datasets, improving its multi-view consistency and 3D geometry. Similarly, DreamGaussian utilizes Zero123 to generate multi-view pseudo-GT images, training a Gaussian-based 3D representation, resulting in a better one-image 3D generation. However, Zero123 is fine-tuned on limited-scale 3D datasets, inevitably reducing its generation capability and resulting in limited generation quality. (3) 2D SDS and 3D SDS combined 3D generation. Intuitively, Magic123 [16] proposes a 3D generation method that combines the generative capabilities of SD with the multi-view consistency of Zero123, achieving impressive results. However, it still inherits the limitations of SD (e.g., the Janus problem and content distortion) and requires a time-consuming iterative textual inversion to fit the learnable text prompt to the given image, significantly restricting the generative capacity of this text-to-image (T2I) model. Therefore, our motivation is raised:
How to avoid the limitations of using 2D SD, a text-to-image model, for image-to-3D generation and fully unleash its generative capability?
Insight. High-quality 3D generation requires three conditions: multi-view consistent geometry, clear textural details, and content that matches the input image. (1) Spatial Domain: As shown in Fig. 2 (a), we visualize the generation results of SD ($S_{2D}$) and Zero123 ($S_{3D}$) in the spatial (RGB) domain. Although the generated image by Zero123 ensures reasonable 3D structure and content matching with the input, it is overly smooth and lacks details. Conversely, the generated image by SD exhibits clear textures but alters content, leading to mismatches with the input image. Therefore, combining SD and Zero123 in the spatial domain is not the optimal solution. (2) Frequency Domain: We further transform the images to the frequency domain using Discrete Fourier Transform (DFT), as shown in Fig. 2 (b). The result of SD ($F_{2D}$) exhibits more mid-to-high frequency components (frequency increases outwards) compared to the result of Zero123 ($F_{3D}$). According to the principle of Fourier transform, higher frequencies represent finer textures. Consequently, the generation result of SD maintains its advantage of detailed textures in the frequency domain, and since it is not in the spatial domain, it does not forcibly constrain the RGB content, but only the degree of texture details. Based on the above analysis, we propose our perspective on one-image 3D generation:
In the spatial domain, use Zero123 to ensure reasonable 3D geometry and content that matches the input. In the frequency domain, use SD to enrich the texture details.
This solution effectively addresses the two motivations we outlined above, achieving high-quality one-image 3D generation. (1) Avoiding the limitations of SD: Since the Janus problem and content distortions arise in the spatial domain, we avoid using SD in the spatial domain. (2) Unleashing the generative power of SD: Similarly, by avoiding strong supervision on the input content in the spatial domain, we can discard the textual inversion and a high CFG value, avoiding SD fitting to local solutions and limiting the generation capability.
@@
    We use both spatial and frequency supervision for image-to-3D. The proposed hybrid Fourier Score Distillation (hy-FSD) fully unleashes the generation ability of the text-to-image model while mitigating the Janus problem and integrates its generation capabilities with the 3D priors of the novel view generation model.
@@
    We develop an efficient image-to-3D generation pipeline, Fourier123, utilizing hy-FSD. It enables high-quality 3D generation in one minute on a single NVIDIA 4090 GPU.
@@
    Extensive experiments confirm that our method significantly enhances the performance of existing optimization-based 3D generation methods, effectively produces 3D assets with reliable structure and elegant appearance.
@@ 2 Related Work
Recently, various 3D representation techniques have been proposed for a range of 3D tasks. Wang et al. [17] employed volumetric rendering and reconstructed object surfaces by training an implicit network. Mildenhall et al. [18] further proposed NeRF, an end-to-end model popular for enabling 3D optimization with only 2D supervision. NeRF has inspired numerous subsequent studies, including 3D reconstruction [19; 20; 21; 22; 23] and generation [24; 14; 7; 25], but it consumes excessive time for optimization due to its computationally expensive forward and backward passes. Although some methods [26; 27; 28] attempted to accelerate training, the recent developed 3D Gaussian Splatting (3DGS) [29] achieved real-time rendering with faster training speed, and is considered a viable alternative 3D representation to NeRF. Its efficient differentiable splatting mechanism and representation design enable fast convergence and faithful reconstruction [30; 31; 32]. Recent studies on 3D generation [33; 34] have adopted 3DGS to achieve faster and higher quality generation. In this work, we also employ 3DGS as the representation technique and make the first attempt to realize the optimization-based methods in both spatial and frequency domains, improving generation quality.
@@ 2.2 Image-to-3D Generation
Image-to-3D generation aims to create 3D assets from a single reference image. This problem is also known as single-view 3D reconstruction, but such reconstruction settings [35; 36; 37] are limited by uncertainty modeling and often produce blurry results. Recently, diffusion models [5; 38] have achieved notable success in image generation, including text-to-image (T2I) [39; 8] and novel view synthesis [13; 10; 11; 40; 41]. Several methods have attempted to extend 2D image models for 3D generation [14; 7; 25; 42; 43], but they suffer from long optimization times as they require frequent generation of 2D images to train 3D representations. To address this, some studies explicitly injected camera parameters into 2D diffusion models for zero-shot novel view synthesis [10; 40]. Other methods have tried to build end-to-end large reconstruction models to generate 3D assets with a single forward process [6; 44; 45; 46]. However, the quality of the 3D assets generated by these two methods remains rough. Qian et al. [16] found that the T2I model [8] has impressive 2D generation capabilities, while Zero-1-to-3 [10] tends to generate reliable 3D structures. They used both 2D and 3D priors to generate 3D. However, it still remains serious issues in tedious time costs and limited generation capacity. In this paper, we combine 2D and 3D priors from the spatial and frequency domains, respectively, fully utilizing their respective advantages and enhancing 3D generation quality.
@@ 3 Preliminary
We use 3DGS [29] as 3D representation. 3DGS uses anisotropic Gaussians to represent scenes, defined by a center position $\mathbf{\mu}\in\mathbb{R}^{3}$ and a covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$, which is decomposed into a scaling factor $\mathbf{s}\in\mathbb{R}^{3}$ and a rotation factor $\mathbf{r}\in\mathbb{R}^{4}$. Additionally, the color of each 3D Gaussian is defined by spherical harmonic (SH) coefficients $\mathbf{h}\in\mathbb{R}^{3\times(k+1)^{2}}$ for order $k$, along with an opacity value $\sigma\in\mathbb{R}$. The 3D Gaussian can be queried as:
@@
where $\mathbf{x}$ represents the position of the query point. To compute the color of each pixel, it uses a typical neural point-based rendering [47]. Let $\mathbf{C}\in\mathbb{R}^{H\times W\times 3}$ represent the color of rendered image, where $H$ and $W$ are the height and width. Rendering process is outlined as:
@@
where $\alpha_{t}\in(0,1)$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. $\bm{\epsilon}_{\theta}$ is trained by minimizing the noise reconstruction loss conditioned on $\mathbf{y}$ from pre-trained language models [48]:
@@
where $\mathcal{C}(\mathbf{R},\mathbf{T})$ is the embedding of the input view and camera extrinsics. After training, users input the image $\mathbf{x}$ and camera external parameters $\mathbf{R},\mathbf{T}$, obtaining the target view in the appropriate pose.
@@
Score Distillation Sampling (SDS) [14] utilizes pre-trained text-to-image diffusion models to optimize the parameters $\bm{\phi}$ of a differentiable 3D representation, such as a neural radiance field or 3DGS. The loss gradient $\mathcal{L}_{\textrm{SDS}}$ is:
@@
where $\mathbf{x}=g(\bm{\phi},c)$ represents an image rendered from the 3D representation $\bm{\phi}$ by the renderer $g$ under a specific camera pose $c$. The weighting function $w(t)$ depends on the timestep $t$, and the noise $\bm{\epsilon}$ is added to $\mathbf{z}=\mathcal{E}({\mathbf{x}})$ following Eq. 3 at timestep $t$. Its key insight is to enforce the rendered image of the learnable 3D representation to adhere to the distribution of the pre-trained diffusion model.
@@ 4 Proposed Fourier123
In this section, we first illustrate the proposed novel score distillation sampling function, hy-FSD (hybrid Fourier Score Distillation), that supervises 3D generation in both spatial and frequency domains. hy-FSD enables the produced 3D assets to benefit from the high-quality generation capability of a text-to-image (T2I) model and the geometric prior of a novel view generation diffusion model simultaneously. Experiments validate that hy-FSD can be applied to existing 3D generation baselines for performance gains. Next, we present an efficient image-to-3D generation pipeline, Fourier123. It generates more reliable geometric structures and high-quality appearances.
@@ 4.1 Hybrid Fourier Score Distillation
As mentioned in Sec. 1, to fully utilize leading generation ability of SD whilst avoiding its natural limitations, we propose to use SD in frequency domain for enriching texture details of 3D assets with Fourier transform. Meanwhile, we adopt RGB results of Zero123 to ensure reasonable 3D geometry.
4.1.1 Fourier Score for Stable Diffusion
The Discrete Fourier Transform (DFT), noted $\mathcal{D}$, has been widely used to analyze the frequency components of images. For multichannel color images, $\mathcal{D}$ is computed and applied independently to each channel. For simplicity, here we omit the notation related to channels. For an image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, $\mathcal{D}$ converts it to the frequency domain as the complex component $\mathbf{X}$, expressed as:
$\mathcal{D}(\mathbf{x})_{u,v}=\mathbf{X}_{u,v}=\frac{1}{\sqrt{HW}}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\mathbf{x}_{h,w}e^{-j2\pi(u\frac{h}{H}+v\frac{w}{W})}.$ (7)
This can be efficiently implemented with the FFT algorithm described in [49]. Note that $\mathbf{X}$ contains phase and amplitude components, and the latter $\mathcal{A}(\mathbf{x})_{u,v}$ is formulated as: